# Uncovering Neuroimaging Biomarkers of Brain Tumor Surgery with AI-Driven Methods

### v1

@@ 4.2 Ablation Study of Unsupervised Pretraining and Fine-Tuning Strategies
We conducted an ablation study comparing two different cohort training strategies (see 2.2.1). Based on Table 2, the best validation results were achieved using the second cohort strategy, particularly with the Addenbrooke’s Hospital and BraTS2025 datasets. Specifically, the Swin-Unet model achieved the lowest error values across both strategies, with the best performance in the second strategy; an RMSE of 0.008 compared to 0.010, MSM of 0.001 in both cases, and MAE of 0.005 compared to 0.009. These results highlight the superiority of the strategy involving both the Addenbrooke’s Hospital and BraTS2025 datasets over the strategy using only the Addenbrooke’s Hospital dataset.
@@
The performance outcomes for fine-tuning in the survival binary classification task using sMRI data (see 2.2.2) are illustrated in Figure 4. We conducted an ablation study across three encoder-decoder configurations: Swin-Unet with an MLP output layer, Swin-Unet with an attention-based output layer, and a baseline AutoEncoder. Each model was evaluated under two fine-tuning strategies: (i) freezing the pre-trained encoder, and (ii) unfreezing the encoder during downstream training. The variability reported reflects results obtained via 5-fold CV. As shown in Figure 4a, frozen encoders exhibited higher variability across most metrics compared to their unfrozen counterparts. Surprisingly, the frozen configurations also achieved higher average performance. Among all models, the Swin-Unet with an attention-based output layer and frozen encoder achieved the best overall results, with an average F1-score of 0.52, accuracy of 0.67, sensitivity of 0.64, and precision of 0.55. Its maximum values across folds reached an F1-score of 0.56, accuracy of 0.77, sensitivity of 0.66, and precision of 0.65. Although the AutoEncoder architecture achieved slightly higher maximum values in F1-score and sensitivity, it exhibited considerably higher CV variability across all metrics and consistently lower precision (below 0.57), indicating a higher false-positive rate compared to the Swin-Unet with the attention-based output layer. These findings highlight a trade-off between performance stability and sensitivity, and suggest that attention-based decoding in transformer-style architectures offers a more reliable and interpretable solution for domain-specific fine-tuning in neuroimaging applications.
Lastly, Figure 4b provides further evidence from the ablation study on different unsupervised training cohorts in the fine-tuning task, confirming the superiority of the strategy that leverages both the Addenbrooke’s Hospital and BraTS2025 datasets compared to the approach that relies solely on the Addenbrooke’s Hospital dataset. Training from scratch on the Addenbrooke’s Hospital dataset without any fine-tuning resulted in substantially poorer performance compared to either of the two unsupervised training cohorts strategies.
@@ 4.3.1 Metrics and interpretations of XAI models
The proposed global explanation optimizer outperformed both the baseline explanation methods used during its training and testing; namely, Gradient SHAP, Guided Backpropagation, and Guided Grad-CAM, as well as established explanation techniques not involved in its training process, including Input × Gradient [21], Integrated Gradients[21], and Kernel SHAP [20]. In terms of faithfulness, the optimizer achieved a score of 0.913 (see Table 3). It also had the lowest average RMSE (0.964), MAE (0.610), and MSM (0.967). Standard deviation was assessed across four global explanations: pre- and post-surgery as well as shorter-term and longer-term survivals. While Guided GradCam showed the highest sparseness (0.612), its faithfulness was below 0.362. The optimized method had the highest reliability aligning closely with the first PCA component of sMRI images and preserving key PCA-derived features. Figure 2(b) illustrates results for the post-surgery longer-term survivals and pre-surgery shorter-term survivals using different $l_{1}$, $l_{2}$, and $l_{3}$ parameter combinations from (4).
@@ 4.3.2 Patterns identified in XAI explanations
Figure 5 displays the Euclidean distances between pre- and post-surgery scans for both the longer-term and shorter-term survival cohorts. These distances are shown for the first PCA component derived from local explanations, the local explanations themselves (Gradient SHAP, Guided Backpropagation, and Guided GradCAM), as well as for the global explanation (Global explanation). The global explanation optimizer outperforms the other methods in terms of sparsity and faithfulness, offering better insights into the global patterns. Thus, we focus primarily on discussing the results from the first PCA component obtained from local explanations and the global explanation maps in Figure 5. By comparing with the atlas using the same thresholding criterion as applied in the PCA Euclidean distance maps, at least 50% of voxels exceeding the 80th percentile and at least one voxel above the 95th percentile, the significant regions are summarised in Table 1 (columns First PCA from local explanations and Global optimizer explanation).
@@
Overlap patterns between both (first PCA and global optimizer) explanations show that for the longer-term survival cohort, the common regions are Orbital and Polar Frontal, while for the dead cohort, the overlapping regions are Auditory Association (AA), Medial Temporal (MT) and Orbital and Polar Frontal (OPF). The superiority of the global explanation optimizer results is evident, as it includes regions that are observed in the Euclidean distance to the entire PCA component space (see 4.1). For example, the Posterior Cingulate (PC) for the shorter-term survival cohort and the Anterior Cingulate and Medial Prefrontal (ACMP) regions for the longer-term survival cohort are identified. These regions are not included in the first PCA component highlighted in Figure 5.
@@ 5 Discussion
A suggested guidance based on Table 1 follows the pattern below: Feature engineering (PCA-based Euclidean and cosine maps) revealed that the surgery regions, i.e. those showing the largest pre- vs. post-surgery changes within tumour masks PCA-space, partially overlap with (and may help explain) the post-surgery alterations observed in brain regions. A key example is the Early Auditory (EA) cortex, which consistently appeared across both survival groups and map types in both the surgical and brain-level results, suggesting it is a core region affected by tumour resection and a hub of post-operative reorganisation [32]. Similarly, Insular and Frontal Opercular (IFO) areas and the OPF cortex were commonly involved, indicating that disruption to sensory and frontal integration areas may play a central role in shaping global connectivity changes [33]. In longer-term survivors, surgical effects were more confined to frontal and midline structures (e.g. ACMP), with downstream changes in executive and motor regions, possibly engaging compensatory networks such as the frontoparietal control system. In contrast, shorter-term survivors showed surgical involvement in posterior and multimodal sensory areas (e.g. Posterior Opercular, PO, or Ventral Stream Visual, VSV), paralleled by more diffuse alterations in visual and perceptual cortices, which may reflect greater network fragility or reduced plasticity.
The last two columns of Table 1, representing local and global explanation methods, highlight key brain regions contributing to group distinctions in the binary severity classification. In longer-term survivals, local explanations highlight frontal and temporal areas such as Inferior Frontal (IF), Lateral Temporal (LT) or MT regions. These areas support language, memory, executive functions, and motor planning, indicative of preserved or adaptable networks facilitating recovery. Global explanations in this group further stress integrative hubs like the ACMP, Dorsolateral Prefrontal Cortex (DLP), MT, or OPF, which are implicated in emotional regulation, high-order cognition, and multisensory integration [34]. In shorter-term survivors, the first PCA component based on local explanations also include frontal regions and MTV but is more focused on posterior sensory and association cortices, such as the Dorsal Stream Visual (DSV) and VSV areas. This suggests a heavier impact on visual and interoceptive processing systems that may have less neuroplastic potential.
Overall, the patterns observed across PCA-derived feature maps and model explanation methods suggest that focal surgical changes, particularly in frontal and opercular regions such as EA, IFO, and OPF, are associated with broader post-operative alterations in structurally and functionally connected brain areas. Longer-term survivors consistently show more engagement of fronto-cingulate and temporal structures (e.g., DLP, ACMP, PLMC), linked to executive function and cognitive control, whereas shorter-term survivors exhibit more extensive involvement of posterior sensory and visual association regions (e.g., PO, DSV, VSV), suggesting less focal and potentially less compensable network disruptions [35]. Figure 6 illustrates the regions with the most significant differences in both severity-related groups. The recurrence of regions such as EA and OPF across all analysis methods and survival groups highlights their centrality in the brain’s structural and functional adaptation to tumour resection. These findings support the notion that the surgical impact on specific cortical hubs may influence the extent and efficacy of post-operative neuroplasticity, ultimately modulating clinical outcomes.
A major limitation of this study lies in the limited availability of paired pre- and post-surgical structural MRI data, which restricts the statistical power and the generalisability of the results. Longitudinal imaging of brain tumours remains scarce in clinical datasets, particularly when considering real-world cases with pre- and post-surgical scans. While large public datasets are increasingly available, they often lack longitudinal follow-up or rely on synthetic or preprocessed data that may not fully reflect clinical variability. In contrast, our dataset is composed of real clinical cases, collected under routine care conditions. We believe this provides added value, as it captures the heterogeneity, imaging artefacts, and surgical effects that are often absent in curated public repositories. This specificity allows us to more accurately model the structural impact of surgery on brain tissue and better capture individual variability in post-operative brain reorganisation.
To address this sample size limitation, we are actively collecting additional longitudinal cases to increase cohort size and improve model reliability. In parallel, future work will focus on applying the current framework to other neuroimaging modalities, such as functional MRI (fMRI) or diffusion MRI (dMRI). These modalities offer complementary information that could enhance both the predictive performance of survival models and the interpretability of regional alterations. Integrating functional and structural perspectives will allow a more comprehensive understanding of surgical impact and disease progression.
@@ 6 Conclusions
Our proposed framework integrates XAI with neuroimaging-based feature engineering to predict the survival of brain tumor patients, providing guidance for surgical decision-making to achieve the necessary onco-functional balance. Our findings correlate dissimilarities between tumor volumes and surgical removal areas with their structural impact on the brain post-operation. By extracting global explanations based on DL approaches for predicting short- and long-term survival, the framework serves as a predictive guideline. The results highlight the involvement of sensory and cognitive regions, with greater disruptions observed in shorter-term survivors. The proposed global explanation optimizer enhances biomarker identification by improving faithfulness and interpretability compared to alternative global XAI methods, while also mitigating inter-method global explanation variability that undermines the trustworthiness of explainable AI. It effectively distinguishes between groups, emphasizing its role in identifying survival-related variability.
@@ Acknowledgment
This project has received funding from the European Union’s Horizon Europe research and innovation program under grant agreement No 101147319 (EBRAINS 2.0). It was also supported in part by the PID2022-137629OA-I00 and PID2022-137451OB-I00 projects, funded by the MICIU/AEI/10.13039/and by “ERDF/EU”. C.J.M is supported by grant JDC2023-051807-I funded by MICIU/AEI/10.13039/501100011033 and by ESF+. All research at the Department of Psychiatry in the University of Cambridge is supported by the NIHR Cambridge Biomedical Research Centre (NIHR203312) and the NIHR Applied Research Collaboration East of England. The views expressed are those of the author(s) and not necessarily those of the NIHR or the Department of Health and Social Care.

### v2

@@ 5.2 Ablation Study of Unsupervised Pretraining and Fine-Tuning Strategies
We conducted an ablation study comparing two different cohort training strategies (see 3.2.1). Based on Table 2, the best validation results were achieved using the second cohort strategy, particularly with the Addenbrooke’s Hospital and BraTS2025 datasets. Specifically, the Swin-Unet model achieved the lowest error values across both strategies, with the best performance in the second strategy; an RMSE of 0.008 compared to 0.010, MSM of 0.001 in both cases, and MAE of 0.005 compared to 0.009. These results highlight the superiority of the strategy involving both the Addenbrooke’s Hospital and BraTS2025 datasets over the strategy using only the Addenbrooke’s Hospital dataset.
@@
The performance outcomes for fine-tuning in the survival binary classification task using sMRI data (see 3.2.2) are illustrated in Fig. 4. We conducted an ablation study across three encoder-decoder configurations: Swin-Unet with an MLP output layer, Swin-Unet with an attention-based output layer, and a baseline AutoEncoder. Each model was evaluated under two fine-tuning strategies: (i) freezing the pre-trained encoder, and (ii) unfreezing the encoder during downstream training. The variability reported reflects results obtained via 5-fold CV. As shown in Fig. 4a, frozen encoders exhibited higher variability across most metrics compared to their unfrozen counterparts. Surprisingly, the frozen configurations also achieved higher average performance. Among all models, the Swin-Unet with an attention-based output layer and frozen encoder achieved the best overall results, with an average F1-score of 0.52, accuracy of 0.67, sensitivity of 0.64, and precision of 0.55. Its maximum values across folds reached an F1-score of 0.56, accuracy of 0.77, sensitivity of 0.66, and precision of 0.65. Although the AutoEncoder architecture achieved slightly higher maximum values in F1-score and sensitivity, it exhibited considerably higher CV variability across all metrics and consistently lower precision (below 0.57), indicating a higher false-positive rate compared to the Swin-Unet with the attention-based output layer. These findings highlight a trade-off between performance stability and sensitivity, and suggest that attention-based decoding in transformer-style architectures offers a more reliable and interpretable solution for domain-specific fine-tuning in neuroimaging applications.
Lastly, Fig. 4b provides further evidence from the ablation study on different unsupervised training cohorts in the fine-tuning task, confirming the superiority of the strategy that leverages both the Addenbrooke’s Hospital and BraTS2025 datasets compared to the approach that relies solely on the Addenbrooke’s Hospital dataset. Training from scratch on the Addenbrooke’s Hospital dataset without any fine-tuning resulted in substantially poorer performance compared to either of the two unsupervised training cohorts strategies.
@@ 5.3.1 Metrics and interpretations of XAI models
The proposed global explanation optimizer outperformed both the baseline explanation methods used during its training and testing; namely, Gradient SHAP, Guided Backpropagation, and Guided Grad-CAM, as well as established explanation techniques not involved in its training process, including Input × Gradient [26], Integrated Gradients[26], and Kernel SHAP [25]. In terms of faithfulness, the optimizer achieved a score of 0.913 (see Table 3). It also had the lowest average RMSE (0.964), MAE (0.610), and MSM (0.967). Standard deviation was assessed across four global explanations: pre- and post-surgery as well as shorter-term and longer-term survivals. While Guided GradCam showed the highest sparseness (0.612), its faithfulness was below 0.362. The optimized method had the highest reliability aligning closely with the first PCA component of sMRI images and preserving key PCA-derived features. Fig.2(b) illustrates results for the post-surgery longer-term survivals and pre-surgery shorter-term survivals using different $l_{1}$, $l_{2}$, and $l_{3}$ parameter combinations from (4).
@@ 5.3.2 Patterns identified in XAI explanations
Fig. 5 displays the Euclidean distances between pre- and post-surgery scans for both the longer-term and shorter-term survival cohorts. These distances are shown for the first PCA component derived from local explanations, the local explanations themselves (Gradient SHAP, Guided Backpropagation, and Guided GradCAM), as well as for the global explanation (Global explanation). The global explanation optimizer outperforms the other methods in terms of sparsity and faithfulness, offering better insights into the global patterns. Thus, we focus primarily on discussing the results from the first PCA component obtained from local explanations and the global explanation maps in Fig. 5. By comparing with the atlas using the same thresholding criterion as applied in the PCA Euclidean distance maps, at least 50% of voxels exceeding the 80th percentile and at least one voxel above the 95th percentile, the significant regions are summarised in Table 1 (columns First PCA from local explanations and Global optimizer explanation).
A suggested guidance based on Table 1 follows the pattern below: Feature engineering (PCA-based Euclidean and cosine maps) revealed that the surgery regions, i.e. those showing the largest pre- vs. post-surgery changes within tumour masks PCA-space, partially overlap with (and may help explain) the post-surgery alterations observed in brain regions. A key example is the Early Auditory (EA) cortex, which consistently appeared across both survival groups and map types in both the surgical and brain-level results, suggesting it is a core region affected by tumour resection and a hub of post-operative reorganisation [37]. Similarly, Insular and Frontal Opercular (IFO) areas and the Orbital Polar Frontal (OPF) cortex were commonly involved, indicating that disruption to sensory and frontal integration areas may play a central role in shaping global connectivity changes [38]. In longer-term survivors, surgical effects were more confined to frontal and midline structures (e.g. Anterior Cingulate and Medial Prefrontal, ACMP), with downstream changes in executive and motor regions, possibly engaging compensatory networks such as the frontoparietal control system. In contrast, shorter-term survivors showed surgical involvement in posterior and multimodal sensory areas (e.g. Posterior Opercular, PO, or Ventral Stream Visual, VSV), paralleled by more diffuse alterations in visual and perceptual cortices, which may reflect greater network fragility or reduced plasticity.
The final two columns of Table 1, representing local and global explanation methods, further highlight the regions most relevant for binary survival classification. Among longer-term survivors, local explanations emphasized frontal and temporal areas such as the Inferior Frontal (IF), Lateral Temporal (LT), and Medial Temporal (MT) regions [39]. These regions support language, memory, executive functions, and motor planning, consistent with preserved or adaptable networks facilitating recovery. Global explanations in the same group additionally highlighted integrative hubs such as ACMP, Dorsolateral Prefrontal Cortex (DLP), MT, and OPF, which are associated with emotional regulation, high-order cognition, and multisensory integration [40]. In shorter-term survivors, the first PCA component from local explanations also included frontal and MT regions, but greater emphasis was placed on posterior sensory and association cortices such as the Dorsal Stream Visual (DSV) and VSV, suggesting stronger disruption of visual and interoceptive systems that may be less amenable to functional compensation.
Taken together, the patterns identified across PCA-derived feature maps and model explanation methods suggest that focal surgical changes, particularly in regions such as EA, IFO, and OPF—are linked to broader alterations in structurally and functionally connected brain areas. Longer-term survivors showed more consistent engagement of fronto-cingulate and temporal regions (e.g., DLP, ACMP, PLMC), associated with executive and cognitive control functions, while shorter-term survivors exhibited widespread posterior sensory and visual involvement (e.g., PO, DSV, VSV), consistent with less focal and less compensable network disruption [41]. Fig. 6 illustrates the regions most consistently differentiated between survival groups. The recurrence of hubs such as EA and OPF across multiple analytical methods underscores their centrality in post-operative adaptation and suggests that surgical impact on specific cortical hubs may shape the extent of functional reorganisation, thereby influencing clinical outcomes.
@@ 6 Discussion
In this study, we studied how structural brain reorganization after glioma surgery relates to patient survival, leveraging a uniquely curated and rare dataset of paired pre- and post-surgical MRI scans. By integrating XAI with latent-space feature engineering (PCA), we identified survival-related neuroimaging biomarkers and generated cohort-level explanations that improved both interpretability and reliability compared to existing state-of-the-art methods.
Our findings demonstrate that regions undergoing the greatest surgical changes often overlapped with broader post-surgical alterations in brain networks, underscoring the interplay between local resection effects and global connectivity reorganization. A consistent involvement of the Early Auditory (EA) cortex across both survival groups suggests that it may represent a central hub of post-operative plasticity as well as a point of vulnerability [42]. Other regions, including the Insular, Frontal Opercular, and OPF cortices, were repeatedly implicated, highlighting the key role of sensory–frontal integration areas in shaping recovery trajectories [43, 44]. Clear survival-related distinctions emerged. Longer-term survivors exhibited more localized surgical effects in frontal and midline structures, with downstream engagement of executive and motor networks, possibly reflecting the recruitment of compensatory systems such as the frontoparietal control network. In contrast, shorter-term survivors displayed greater involvement of posterior and multimodal sensory areas, together with diffuse alterations in visual and perceptual cortices. These contrasting profiles suggest differences in network resilience and neuroplasticity, pointing to potential imaging markers for surgical planning, risk stratification, and post-operative rehabilitation strategies [45]. Overlap patterns between explanatory methods further support these distinctions. For the longer-term survival cohort, both the first PCA component and the global optimizer highlighted Orbital and Polar Frontal regions, whereas in the shorter-term survival cohort, overlap was observed in the Auditory Association, MT, and OPF regions. Importantly, the superiority of the global explanation optimizer is evident in its ability to identify additional survival-related hubs, such as the Posterior Cingulate in shorter-term survivors and the ACMP regions in longer-term survivors, which were not captured by the first PCA component alone (see Section 5.1, Fig. 5). These findings illustrate how optimized global explanations can provide more consistent and clinically meaningful insights compared to conventional approaches.
The main limitation of this work lies in the restricted availability of paired pre- and post-surgical structural MRI data. Such longitudinal imaging remains extremely scarce in clinical practice due to the clinical, logistical, and ethical challenges of acquisition. Although large public datasets are increasingly accessible, they typically lack longitudinal follow-up or rely on synthetic or heavily preprocessed data, which may not adequately reflect clinical variability. By contrast, our dataset consists of real-world clinical cases collected under routine care, capturing the heterogeneity, imaging artefacts, and surgical effects that are often absent in curated repositories. This rarity constitutes a unique strength of the study, enabling us to more accurately model the structural consequences of surgery and better capture individual variability in post-operative brain reorganization. To mitigate the limitations of sample size, we are actively expanding the collection of longitudinal cases to strengthen statistical power and reproducibility. Future work will also extend the framework to additional neuroimaging modalities, including functional and diffusion MRI and diffusion MRI. These complementary modalities can enhance predictive performance, enrich interpretability, and provide a more comprehensive view of both structural and functional brain dynamics. Integrating multimodal imaging perspectives will ultimately advance our understanding of surgical impact, recovery mechanisms, and disease progression, and further reinforce the translational potential of XAI-driven neuroimaging in precision neuro-oncology.
@@ 7 Conclusions
Our proposed framework integrates XAI with neuroimaging-based feature engineering to predict survival in brain tumor patients, offering guidance for surgical decision-making to achieve the critical onco-functional balance. A unique strength of this work lies in the use of a rare, clinically collected dataset comprising paired pre- and post-surgical MRI scans—data that are exceptionally scarce due to the clinical complexities, operational demands, and ethical considerations of acquiring longitudinal datasets. Using this unique dataset, we demonstrate how dissimilarities between tumor volumes and surgical resection areas correlate with their structural impact on the brain post-operatively. By extracting global explanations from deep learning models for predicting short- and long-term survival, the framework functions as a clinically relevant predictive guideline. Our results highlight the consistent involvement of sensory and cognitive regions, with greater disruptions observed in shorter-term survivors, underscoring the importance of preserving networks critical for cognition and perception. Methodologically, the proposed global explanation optimizer improves both faithfulness and interpretability compared to alternative global XAI methods, while reducing inter-method variability that often undermines the trustworthiness of explainable AI. Overall, this work not only establishes a novel XAI-based framework for survival assessment but also demonstrates the scientific and clinical value of rare pre- and post-surgical datasets in uncovering survival-related variability, ultimately advancing precision medicine in neuro-oncology.
@@ Acknowledgment
This project has received funding from the European Union’s Horizon Europe research and innovation program under grant agreement No 101147319 (EBRAINS 2.0). It was also supported in part by the PID2022-137629OA-I00 and PID2022-137451OB-I00 projects, funded by the MICIU/AEI/10.13039/and by “ERDF/EU”. This study was funded from the National Institute for Health and Care Research (NIHR), Career Development Fellowship (CDF-2018-11-ST2-003 to SJP.). This publication presents independent research funded by the National Institute for Health and Care Research (NIHR). All research at the Department of Psychiatry in the University of Cambridge is supported by the NIHR Cambridge Biomedical Research Centre (NIHR203312) and the NIHR Applied Research Collaboration East of England. The views expressed are those of the author(s) and not necessarily those of the NHS, the NIHR or the Department of Health and Social Care. C.J.M is supported by grant JDC2023-051807-I funded by MICIU/AEI/10.13039/501100011033 and by ESF+.